OpenStudy (anonymous):

If on a scale drawing 5 inches represents 80 feet, then inch would represent how many feet? 16 feet 12 feet 8 feet 4 feet

OpenStudy (anonymous):

16feet

OpenStudy (anonymous):

5x=80 x=80/5=16

OpenStudy (lalaly):

inches feet 5 80 1 x cross multiply 5x=80 x=80/5 =16 feet

OpenStudy (ledah):

...I think they would've benefited from the explanation first :c

OpenStudy (anonymous):

thanx everyone